# Smooth polynomials with several prescribed coefficients

### v2

@@ 1. Introduction
For a given index-set ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ with #$\#{\mathcal{I}}\leq\delta(g)\cdot n$ and $0\in{\mathcal{I}}$, it follows from [2, 25], that the number of prime numbers with prescribed digits $\alpha_{i}\in\{0,1,\dots,g-1\}$ ($i\in{\mathcal{I}}$) with $\alpha_{0}\neq 0$ satisfies
#$\#\left\{p<g^{n},p_{i}=\alpha_{i},i\in{\mathcal{I}},\text{$p$ prime}\right\}\sim\frac{\pi(g^{n})}{g^{\#{\mathcal{I}}}},$
@@
#$\#({\mathcal{P}}(n)\cap{\mathcal{J}})=\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}\left(1+O\left(\frac{1-\log_{q}\delta}{q^{1/\delta-4/(\delta+1)}}\right)\right)+O(q^{3n/4}),$
@@
In particular, we have the expected frequency of irreducible polynomials with prescribed coefficients, #$\#({\mathcal{P}}(n)\cap{\mathcal{J}})\sim\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}$, if $q\rightarrow\infty$ or $\delta=\#{\mathcal{I}}/n\rightarrow 0$.
@@
holds in range $n\geq m\geq\sqrt{n\log n}$ (see also [4, 8, 23]). Recently, Gorodetsky [11] gave better approximation of $\Psi(n,m)$ with main term $q^{n}\rho(u-n/(2m^{2}))$ and improved error term.
@@
Let $40<\varepsilon<1/$ 4. Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$) such that $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$. Write $\delta=\#{\mathcal{I}}/n$ and assume that
$\delta<1/10,\quad\text{that is,}\quad\#{\mathcal{I}}<n/10.$
Then we have
(4) #$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#I}}\left(1+O\left(\left(u\log u\right)^{1/8}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}+\frac{1}{q}\frac{u^{1/2}(\log u)^{5/2}}{m}\right)\right)\\ &\qquad+O\left(\frac{q^{n}}{q^{\#I}}q^{-\frac{(1-\varepsilon)n+2m}{8}+2}+m^{2}n^{1/2}q^{(7n+m+4)/8}\right),\end{split}$
holds for some positive $C>0$ in range $n\geq m\geq\sqrt{n\log n}$, where the implied constant may depend on $\varepsilon$.
@@
Assume, that
@@
  • for fixed $q$, we have $n\rightarrow\infty$ and $\delta=o(n)$.
@@
(5) #$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})\sim q^{-\#I}\Psi(n,m).$
If $\alpha_{0}=0$, then #$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})$ deviates from the expected frequency. For example, if $I=\{0,1,\dots,r-1\}$ with $\alpha_{i}=0$ for $i\in{\mathcal{I}}$, then
@@
(6) #$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})=\Psi(n-r,m)\sim\rho\left(\frac{n-r}{m}\right)q^{n-r}$
@@
$I=\{i_{0}<\dots<i_{\nu}<\dots\}$
with $\alpha_{i_{0}}=\dots=\alpha_{i_{\nu-1}}=0\neq\alpha_{i_{\nu}}$ and define
$\Lambda_{0}(n,m,I,\nu)=\frac{-1}{m}\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<m/8\end{subarray}}q^{j-i_{j}}\rho^{\prime}\left(\frac{n-i_{j}}{m}\right)$
and
$\Lambda_{1}(n,m,I,\nu)=\left\{\begin{array}[]{cl}-\frac{1}{m}\frac{q^{\nu-i_{\nu}}}{q-1}\rho^{\prime}\left(\frac{n-i_{\nu}}{m}\right)&\text{if $i_{\nu}<m/8$},\\ 0&\text{otherwise}.\end{array}\right.$
@@
Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and choose $\alpha_{i}\in\mathbb{F}_{q}$ for each $i\in{\mathcal{I}}$. Write $\delta=\#{\mathcal{I}}/n$ and assume that
$\delta<1/10,\quad\text{that is,}\quad\#{\mathcal{I}}<n/10,$
then we have
#$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})$ $=\frac{\Psi(n,m)}{q^{\#I}}\left(1+O\left(\left(u\log u\right)^{1/8}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}+\frac{1}{q^{i_{0}+1}}\frac{u^{1/2}(\log u)^{5/2}}{m}\right)\right)$
$\qquad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)+O\left(q^{-\frac{(1-\varepsilon)n+2m}{8}+2}e^{\varepsilon(n-m)}\right)\right)$
$\qquad+O\left(m^{2}n^{1/2}q^{(7n+m+4)/8}\right),$
holds for some positive $C>0$ in range $n\geq m\geq\sqrt{n\log n}$.
@@
We remark, if $0\in{\mathcal{I}}$, then we can apply either Theorem 1 if $\alpha_{0}\neq 0$ or (6) if $\alpha_{0}=0$. In the case $0\not\in{\mathcal{I}}$, Theorem 3 yields the expected frequency if $q\rightarrow\infty$ or $\delta$ is small enough in terms of $u$.
@@
Assume, that $0\not\in{\mathcal{I}}$ and
  • if $q\rightarrow\infty$ for fixed $n$ and $u$; or
  • if $n\rightarrow\infty$ with $\delta=o\left((u\log u)^{9/8}\right)$ and fixed $q$.
Then (5) holds.
Theorems 1 and 3 give asymptotics to the number smooth polynomials with several prescribed coefficients. Recently, Kuttner and Wang [18], using different techniques, have considered the exact number of smooth polynomials of degree $n$ with given trace, that is, when =${\mathcal{I}}=\{n-1\}$.
@@
The minor arc analysis is based on double exponential sums and the good factorisation properties of smooth polynomials, see Section 5.2.
@@ 2. Outline of the proof
with the convention $|0|=0$. One can extend the absolute value ||$\cdot$|| ⋅ | to $\mathbb{F}_{q}(t)$ in the natural way by $|f/g|=q^{\deg f-\deg g}$.
@@
We investigate this integral by the circle method. To do so, let $\kappa,\ell$ be two positive parameters to be fixed later such that
@@
We give approximation of $S(\xi;n,m)$ on major arcs in Section 5.1 using character sum estimates on smooth polynomials introduced in Section 4. The minor arc estimates of $S(\xi;n,m)$ are based on double exponential sums and the good factorisation properties of smooth polynomials, see Section 5.2.
@@ 3.1. Notations
For given functions $F$ and $G$, the notations $F\ll G$, $G\gg F$ and $F=O(G)$ are all equivalent to the statement that the inequality $|F|\leq c|G|$ holds with some constant $c>0$. Throughout the paper, any implied constants in symbols O, $\ll$ and $\gg$ are absolute unless specified otherwise. For a set ${\mathcal{A}}$ we denote the characteristic function of ${\mathcal{A}}$ by $\mathbf{1}({\mathcal{A}})$.
@@ 3.2. Arithmetically distributed relations
Observe, that ${\mathcal{R}}_{\ell}$ on fixed degree polynomials captures the polynomials on intervals, that is, for polynomials $f,h\in{\mathcal{M}}(n)$,
@@
see [15, Theorem 8.6] and [17, Lemma 1.1]. We also have #$\#G_{\ell}=q^{\ell}$ and
#$\#G_{g}=\#(\mathbb{F}_{q}[t]/(g))^{\times}=\Phi(g).$
@@ 3.3. Exponential sums
$\sum_{\chi\bmod{\mathcal{R}}_{\ell,g}}\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a\bmod{\mathcal{R}}_{\ell,g}\\ a\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\chi(a)\mathbf{e}(a/g)\right|^{2}=q^{\ell}\Phi(g)^{2}$
@@
[…]
@@ 3.4. Dickman’s $\rho$ function
For the following result see [7, Lemma 4.1].
Lemma 13.
For any integer $k\geq 0$ we have
$(-1)^{k}\rho^{(k)}(u)>0\quad\text{for $u>1$}$
and
$\rho^{(k)}(u)\sim(-1)^{k}\rho(u)(\log u)^{k}\quad\text{as }u\rightarrow\infty.$
@@ 4. Distribution of smooth polynomials
The purpose of this section is to prove character sum estimates along smooth polynomial. The proof is based on the method of Bhowmick, Lê and Liu [1], see also [10]. We consider both cases of $q$ or $n$ are fixed in contrast to [1, 10].
@@
    $\left|\sum_{f\in S(n,m)}\chi(f)\right|\leq q^{\frac{n}{2}+\frac{\log\log(\ell+1+\deg g)}{\log(\ell+1+\deg g)}}e^{13q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}}$
@@
    Let $40<\varepsilon<1/$ 4, then we have
@@
For for $r\geq R$ we use (15) in (14), while $r<R$ we use the trivial bound (16). Namely, (15) yields that for $r\geq R$, we have that the coefficient of $z^{r}$ in (14) is bounded in absolute value by
@@
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{R(1-q^{-\delta})}\right).$
As $\delta<1/\log q$, we have $1-q^{-\delta}\geq\frac{1}{2}\delta\log q$, thus
$\log T(z,\chi;m)$ $\leq(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{R\delta\log q}\right)$
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right).$
@@
$\leq\exp\left((\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)\right)q^{(1/2+\delta)n}.$
@@
$(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)$
$=$ $\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}\left(13q^{1/2}+\frac{1}{\log\log(\ell+1+\deg g)}\right)$
$\leq$ $13q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}$
@@
Let $\chi\neq\chi_{0}$ be a character modulo ${\mathcal{R}}_{\ell,g}$ and $n\geq\ell+\deg g$. For all $40<\varepsilon<1/$ 4, we have
@@
We can assume, that $q^{n}$ is large enough in terms of $\varepsilon$.
@@
Let $40<\varepsilon<1/$ 4. For $1\leq\ell\leq n$ we have for $b\in\mathbb{F}_{q}[t]$ that
#$\#\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}=\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right),$
@@
(24) #$\#\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}$ $=\frac{1}{q^{\ell}}\sum_{\chi\bmod{\mathcal{R}}_{\ell}}\overline{\chi(b)}\sum_{f\in{\mathcal{S}}(n,m)}\chi(f).$

### v3

@@ 1. Introduction
For a given index-set ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ with ${\mathcal{I}}\leq\delta(g)\cdot n$ and $0\in{\mathcal{I}}$, it follows from [2, 25], that the number of prime numbers with prescribed digits $\alpha_{i}\in\{0,1,\dots,g-1\}$ ($i\in{\mathcal{I}}$) with $\alpha_{0}\neq 0$ satisfies
$\left\{p<g^{n},p_{i}=\alpha_{i},i\in{\mathcal{I}},\text{$p$ prime}\right\}\sim\frac{\pi(g^{n})}{g^{\#{\mathcal{I}}}},$
@@
$({\mathcal{P}}(n)\cap{\mathcal{J}})=\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}\left(1+O\left(\frac{1-\log_{q}\delta}{q^{1/\delta-4/(\delta+1)}}\right)\right)+O(q^{3n/4}),$
@@
In particular, we have the expected frequency of irreducible polynomials with prescribed coefficients, $({\mathcal{P}}(n)\cap{\mathcal{J}})\sim\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}$, if $q\rightarrow\infty$ or $\delta=\#{\mathcal{I}}/n\rightarrow 0$.
@@
holds in range $n\geq m\geq\sqrt{n\log n}$ (see also [4, 8, 23]). Recently, Gorodetsky [11] gave the following better approximation of $\Psi(n,m)$
(4) $\Psi(n,m)=q^{n}\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)$
uniformly for $(2+\varepsilon)\log_{q}n\leq m\leq n$ and
(5) $\Psi(n,m)=q^{n}\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)$
if $\sqrt{n\log n}=O(m)$. Moreover, the error terms in (3) and (5) are optimal.
@@
Let $0<\varepsilon<1/4$. Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$) such that $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$. Assume,
(6) $m\leq n/100,\quad\text{that is,}\quad u\geq 100.$
Write $\delta=\#{\mathcal{I}}/n$ and assume that
$\delta<1/24,\quad\text{that is,}\quad\#{\mathcal{I}}<n/24.$
Then for
(7) $(2+\varepsilon)\log_{q}n\leq m\leq n,$
we have
(8) $\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\right)\right)\\ &\qquad+O\left(\frac{q^{n}}{q^{\#{\mathcal{I}}}}m^{2}n^{1/2}q^{-\eta n}\right),\end{split}$
holds for some positive $C>0$ and $\eta>0$, where the implied constant may depend on $\varepsilon$.
Theorem 1 gives bounds on $({\mathcal{S}}(n,m)\cap{\mathcal{J}})$. However, in a shorter range for $m$ instead of (7) we also have an asymptotic result.
Corollary 2.
Having the assumptions as in Theorem 1, for
(9) $\sqrt{n\log n}=O(m),$
we have
(10) $\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)\right)\\ &\qquad+O\left(\frac{q^{n}}{q^{\#{\mathcal{I}}}}m^{2}n^{1/2}q^{-\eta n}\right),\end{split}$
holds for some positive $C>0$ and $\eta>0$.
@@
Assume that the conditions of Corollary 2 hold. Moreover, if
@@
  • for fixed $q$, we have $n\rightarrow\infty$ and $\delta=o(1)$,
@@
(11) $({\mathcal{S}}(n,m)\cap{\mathcal{J}})\sim q^{-\#{\mathcal{I}}}\Psi(n,m).$
If $\alpha_{0}=0$, then $({\mathcal{S}}(n,m)\cap{\mathcal{J}})$ deviates from the expected frequency. For example, if $I=\{0,1,\dots,r-1\}$ with $\alpha_{i}=0$ for $i\in{\mathcal{I}}$, then
@@
(12) $({\mathcal{S}}(n,m)\cap{\mathcal{J}})=\Psi(n-r,m)\sim\rho\left(\frac{n-r}{m}\right)q^{n-r}$
@@
$I=\{i_{0}<\dots<i_{\nu}<\dots\}\quad\text{with }\quad\alpha_{i_{0}}=\dots=\alpha_{i_{\nu-1}}=0\neq\alpha_{i_{\nu}}$
and for $0<\kappa<n$ define
$\Lambda(n,m,{\mathcal{I}},\kappa)=$ $\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<\kappa\end{subarray}}\frac{1}{q^{n-j-1}}\left(\Psi(n-i_{j}-1,m)-\frac{1}{q}\Psi(n-i_{j},m)\right)$
$-\mathbf{1}(\nu<\kappa)\frac{1}{q^{n-\nu-1}}\left(\frac{1}{q-1}\Psi(n-i_{\nu}-1,m)-\frac{1}{q(q-1)}\Psi(n-i_{\nu},m)\right).$
@@
$0<\varepsilon<1/4$. Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$).
$m\leq n/100,\quad\text{that is,}\quad u\geq 100.$
Write $\delta=\#{\mathcal{I}}/n$ and assume that
$\delta<1/24,\quad\text{that is,}\quad\#{\mathcal{I}}<n/24.$
Then for
(13) $(2+\varepsilon)\log_{q}n\leq m\leq n,$
and
(14) $0<\kappa\leq n/5,$
we have
$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}},\kappa)+O\left(nq^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa+\frac{3}{2}}e^{\varepsilon n}\right)\right)+O\left(m^{2}n^{1/2}q^{n-\kappa/2}\right),\end{split}$
holds for some positive $C>0$, where the implied constant may depend on $\varepsilon$.
Again, if $m$ is in the shorter range (9) instead of (13), a stronger result can be obtained for $({\mathcal{S}}(n,m)\cap{\mathcal{J}})$.
Corollary 5.
Having the assumptions as in Theorem 4, in the range (9) we have
$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}},\kappa)+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa+\frac{3}{2}}e^{\varepsilon\kappa}\right)\right)+O\left(m^{2}n^{1/2}q^{n-\kappa/2}\right),\end{split}$
holds for some positive $C>0$, where the implied constant may depend on $\varepsilon$.
@@
In particular, we remark that if $0\in{\mathcal{I}}$, then we can apply either Theorem 1 if $\alpha_{0}\neq 0$ or (12) if $\alpha_{0}=0$. In the case $0\not\in{\mathcal{I}}$, Theorem 4 yields the expected frequency if $q\rightarrow\infty$.
@@
Assume that the conditions of Theorem 4 hold and $n,m$ satisfy (9). If $0\not\in{\mathcal{I}}$ and $q\rightarrow\infty$ for fixed $n$ and $u$, then (11) holds.
Theorems 1 and 4 give asymptotics to the number smooth polynomials with several prescribed coefficients. Recently, Kuttner and Wang [18], using different techniques, have considered the exact number of smooth polynomials of degree $n$ with given trace, that is, when ${\mathcal{I}}=\{n-1\}$.
@@
The minor arc analysis is based on double exponential sums and the good factorization properties of smooth polynomials, see Section 5.2.
Finally, Theorem 1 follows from Theorem 4, see Section 8.
@@ 2. Outline of the proof
with the convention $|0|=0$. One can extend the absolute value $|\cdot|$ to $\mathbb{F}_{q}(t)$ in the natural way by $|f/g|=q^{\deg f-\deg g}$.
@@
We investigate this integral by the circle method. To do so, let $\ell$ be a positive parameters to be fixed later such that
@@
We give an approximation of $S(\xi;n,m)$ on major arcs in Section 5.1 using character sum estimates on smooth polynomials introduced in Section 4. The minor arc estimates of $S(\xi;n,m)$ are based on double exponential sums and the good factorization properties of smooth polynomials, see Section 5.2.
@@ 3.1. Notations
For given functions $F$ and $G$, the notations $F\ll G$, $G\gg F$ and $F=O(G)$ are all equivalent to the statement that the inequality $|F|\leq c|G|$ holds with some constant $c>0$. Throughout the paper, any implied constants in symbols O, $\ll$ and $\gg$ are absolute, unless specified otherwise. For a set $\mathcal{A}$} we denote the characteristic function of $\mathcal{A}$} by $\mathbf{1}({\mathcal{A}})$.
@@ 3.2. Arithmetically distributed relations
Observe that $\mathcal{R}$$_{\ell}$ on fixed degree polynomials captures the polynomials on intervals, that is, for polynomials $f,h\in{\mathcal{M}}(n)$,
@@
see [15, Theorem 8.6] and [17, Lemma 1.1]. We also have $G_{\ell}=q^{\ell}$ and
$G_{g}=\#(\mathbb{F}_{q}[t]/(g))^{\times}=\Phi(g).$
@@
$\sum_{\chi\bmod{\mathcal{R}}_{\ell,g}}\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a\bmod{\mathcal{R}}_{\ell,g}\\ a\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\chi(a)\mathbf{e}(a/g)\right|^{2}=q^{\ell}\Phi(g)^{2}.$
@@
$=\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a_{1},a_{2}\bmod{\mathcal{R}}_{\ell,g}\\ a_{1}\equiv a_{2}\equiv b\mod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}((a_{1}-a_{2})/g)\sum_{\chi\bmod{\mathcal{R}}_{\ell,g}}\chi(a_{1})\bar{\chi}(a_{2})$
@@ 4. Distribution of smooth polynomials
The purpose of this section is to prove character sum estimates along smooth polynomials. The proof is based on the method of Bhowmick, Lê and Liu [1], see also [10]. We consider both cases that $q$ or $n$ goes to infinity, in contrast to [1, 10].
@@
    $\left|\sum_{f\in S(n,m)}\chi(f)\right|\leq q^{\frac{n}{2}+\frac{\log\log(\ell+1+\deg g)}{\log(\ell+1+\deg g)}}e^{39q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}}$
@@
    Let $0<\varepsilon<1/4$, then we have
@@
For $r\geq R$ we use (23) in (22), while $r<R$ we use the trivial bound (24). Namely, (23) yields that for $r\geq R$, we have that the coefficient of $z^{r}$ in (22) is bounded in absolute value by
@@
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2+\delta}+\frac{1}{R(1-q^{-\delta})}\right).$
As $\delta<1/\log q$, we have $q^{\delta}<3$ and $1-q^{-\delta}\geq\frac{1}{2}\delta\log q$, thus
$\log T(z,\chi;m)$ $\leq(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{R\delta\log q}\right)$
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right).$
@@
$\leq\exp\left((\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)\right)q^{(1/2+\delta)n}.$
@@
$(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)$
$=$ $\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}\left(39q^{1/2}+\frac{1}{\log\log(\ell+1+\deg g)}\right)$
$\leq$ $39q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}$
@@
Let $\chi\neq\chi_{0}$ be a character modulo ${\mathcal{R}}_{\ell,g}$ and $n\geq\ell+\deg g$. For all $0<\varepsilon<1/4$, we have
@@
We can assume that $q^{n}$ is large enough in terms of $\varepsilon$.
@@
Let $0<\varepsilon<1/4$. For $1\leq\ell\leq n$ we have for $b\in\mathbb{F}_{q}[t]$ that
$\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}=\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right),$
@@
$\#\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}$ $=\frac{1}{q^{\ell}}\sum_{\chi\bmod{\mathcal{R}}_{\ell}}\overline{\chi(b)}\sum_{f\in{\mathcal{S}}(n,m)}\chi(f).$